find y(t) solution of the ivp: y''+8y'+20y=-4£(t-2),y(0)=0,y'(0)=1 where £ is the S shaped character
show work
Answer #1

solh j+ @y+20y = 48(+-2) Ylo) 5o ylo)= -25 8 - 4 é (sy-4) +8($g) + roy y ( 5?785 +20) - 472 +1 -25 4 e = ys) (8 +85420) (5